OpenStudy (anonymous):

the numbers on karls three raffle tickets are consecutive integers. The sum of the numbers is 138. What are the numbers?

OpenStudy (anonymous):

1st # = x 2nd # = x+1 What would be the third number?

OpenStudy (anonymous):

138/3 = 46 Therefore... 46 + 46 + 46 = 138 (46-1) + 46 + (46+1) = 138 45 + 46 + 47 = 138.

OpenStudy (anonymous):

my answer is 45, 46 & 47?

OpenStudy (anonymous):

representation: 1 # = x 2# = x+1 3# = x+2 equation: x+(x+1)+(x+2)=138 x+x+1+x+2 =138 3x+3 = 138 3x = 138-3 x = 135/3 x= 45 therefor the numbers are: 45,46,47.